Find the 2nd order parabola (Ŷ = a0 + a1 x + a2 x2) that fits to the following data set. Then calculate correlation coefficient.

x

0

1

2

3

4

5

y

2.1

7.7

13.6

27.2

40.9

61.1
Answer #1

The given data is inputted in EXCEL here and the SCATTERPLOT is used to obtain the scatterplot here. The TRENDLINE function is used to obtain the scatteplot after that and a 2nd degree polynomial is fitted on the above data.

1589672548253_image.png

therefore the equation here is given as:

y = 2.4786 + 2.3593x + 1.8607x^2

The coefficient of determination is obtained here to be: R2 = 0.9985

Therefore correlation coefficient now is computed here as:

r = \sqrt{R^2} = \sqrt{0.9985} = 0.9992

Therefore 0.9992 is the correlation coefficient here.